2.16 Suppose that P(A) = 0.4, P(B) = 0.5 and P(AB) = 0.2. Find the following: a) P(AUB) b) P(AB) e) PIA(AUB) d) PIAU(AB)
Answer #1

a)P(A u B)=P(A)+P(B)-P(AB)=0.4+0.5-0.2=0.7

b)P(A'B)=P(B)-P(AB)=0.5-0.2=0.3

c)P(A'(A u B))=P(A'B)=0.3

d)P(A u (A'B))=P(A u B)=0.7